If the likelihood of an incomplete order form is 0.10, what is the probability that there...

  1. If the likelihood of an incomplete order form is 0.10, what is the probability that there are no more than one incomplete order forms in a sample of four? Assume binomial. Use 4 decimals.
    1. 0.5385
    2. 0.9477
    3. 0.0842
    4. 0.2916
Answer #1

probability that there are no more than one incomplete order forms in a sample of four

=P(X<=1)=P(X=0)+P(X=1)=4C0(0.1)0(0.9)4+4C1(0.1)1(0.9)3=0.9477

option B